OpenStudy (anonymous):

Ana has been recording the amount of times she baby-sat the Chandra children over the past six months. For the first two months, she baby sat for the same amount of time each month. The last two months, she baby-sat 7 times each month. The most she baby-sat them in one month was 10 times and the least amount of times was 2. If the median amount for the six months was 6.5 times, what was the mean number of times Ana baby-sat the Chandra children? Round your answer to the nearest tenth

OpenStudy (yanasidlinskiy):

Hi Jocelyn 1st month = N 2nd month = N 3rd month = 10 4th month = 2 5th month = 7 6th month = 7 (N + 7)/2 = 6.5 N = 13 − 7 [ N = 6 times ] Mean = (1st month + 2nd month + . . . . + 6th month)/6 Mean = (6 + 6 + 10 + 2 + 7 + 7)/6 Mean = 38/6 [ Mean ≃ 6.3 ] Hope this helps !